# Active nematic response to a deformable body or boundary: elastic deformations and anchoring-induced flow

### v1

@@ Abstract
If a boundary is soft, it will generally be deformed by the surface tractions above. We assume the boundary is part of an isotropic solid that deforms according to plane stress/strain in the framework of linear elasticity, with shear modulus $\mu_{s}$ and Poisson’s ratio $\nu$ [72, 73]. In this framework, the deformed (Eulerian) frame and the reference (Lagrangian) frame are mathematically equivalent. At equilbrium, the elastic solid can be described by an Airy-stress function, $\mathcal{A}(x,y)$, which satisfies the biharmonic equation, $\nabla^{4}\mathcal{A}=0$. We can, thus, write
@@
for two (new) Goursat functions, $G$ and $H$, that are locally holomorphic within the solid (but not necessarily the LC), analogous to $g$ and $h$ in (10) [72]. These Goursat functions yield the stress tensor within the solid, $\bm{\Sigma}$, via the Kolosov–Muskhelishvili formulae [66]. Additionally, the solid displacement, $(U,V)$, can be expressed as
@@
up to an arbitrary body displacement and rotation, for the dimensionless elastic modulus $\mathit{M}\coloneqq\mu_{s}L^{2}/K$ and $\kappa=(3-4\nu)$ or $(3-\nu)/(1+\nu)$ for plane strain or plane stress, respectively.
The LC–solid interface must also be at equilibrium, and so, the stresses acting on it must balance: $-\bm{\hat{\nu}}\cdot\bm{\Sigma}=\bm{t}^{e}+\bm{t}^{a}$. Provided we can find functions $G$ and $H$ that are locally holomorphic within the solid and satisfy this stress balance, we will have obtained analytic solutions for the solid’s deformation. However, yet again, the above solution is not invariant under a conformal map, and analytical progress can only be made in specific cases.
We return to the example of a unit cylinder subject to (finite) tangential anchoring. Here, the director angle (11) and streamfunction (12) yield analytic expressions for the elastic traction, $t^{e}_{x}+\mathrm{i}t_{y}^{e}$, and activity-induced traction, $t^{a}_{x}+\mathrm{i}t_{y}^{a}=\mathit{A}\mathrm{Er}(1/z-\rho^{2}z)$, acting on the cylinder. The cylinder is now assumed to be deformable, and so, we seek the Airy stress (14). The Goursat functions $G(z)$ and $H(z)$ must be locally holomorphic in $|z|<1$ and satisfy the stress balance condition on $|z|=1$. See [66] for the solution to this problem.
For passive LCs ($\mathit{A}=0$), the deformation acts to reduce the elastic energy stored in the LC by flattening the cylinder. Figure  2 shows the LC configurations and body deformations for three different anchoring strengths and $\kappa=5/3$ (i.e. an incompressible cylinder, $\nu=1/2$, under plane stress). For large anchoring strengths, $w\gg 1$, most of the LC energy is stored local to the left and right poles of the cylinder, with a singular energy emerging in the case of strong (infinite) anchoring ($w=\infty$) due to the appearance of two $-1$ Boojum topological defects at $z=\pm 1$ [74, 70]. This localization of energy within the LC results in large localized surface tractions: $t^{e}_{x}+\mathrm{i}t_{y}^{e}=\mathcal{O}(w^{2})$ for $z\pm 1=\mathcal{O}(1/w)$, and $t^{e}_{x}+\mathrm{i}t_{y}^{e}=\mathcal{O}(1)$ otherwise, as $w\to\infty$. This, in turn, leads to large localized deformations of the cylinder, see Fig. 2. In particular, we find that
@@
$\mathcal{U}(Z/w)=\kappa/(1+Z)+1/(1+\bar{Z})+(Z+\bar{Z})/(1+\bar{Z})^{2},$ (17)
corresponds to a deformation that is only apparent local to the left and right poles, i.e. when $z\pm 1=\mathcal{O}(1/w)$. As the anchoring strength decreases, stress is no longer concentrated at the poles and the maximum shear stress $\tau(x,y)=\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$ (a common indicator of mechanical failure [73]) is greatest at (±$(\pm x_{\max},0)$, where $x_{\max}$ decreases from $1$ for $w=\infty$, to 00 for $w\leq 24/35$, see Fig. 2.
Figure 2: (top) Deformation of soft cylinders immersed in passive nematic LCs with $\mathit{A}\mathrm{Er}=0$, $\mathit{M}/w=5$, $\kappa=5/3$, and dimensionless anchoring strengths ① $w=0.1$, ② $w=1$, and ③ $w=10$. Color indicates the maximum shear stress, $\tau=\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$, within the cylinder and the blue curves are director lines. (bottom) The positions, $(x_{\max},0)$, and value, $\tau_{\max}$, of the largest maximum shear stress plotted as a function of $w$.
For active LCs ($A\neq 0$), the activity leads to an additional deformation:
$U+\mathrm{i}V=U^{e}+\mathrm{i}V^{e}+\mathit{A}\mathrm{Er}\bigl{[}(1-\kappa)\rho^{2}z+2\bar{z}\bigr{]}/(4\mathit{M}),$ (18)
@@
$\mathcal{A}_{R}\sim 1+U\bigr{|}_{z=1}-V\bigr{|}_{z=\mathrm{i}}=1+(\mathit{A}\mathrm{Er}-\mathcal{W})/\mathit{M},$ (19)
@@
Figure 3: (top) Deformation of soft cylinders immersed in active nematic LCs with $w=5$, $\mathit{M}=50$, $\kappa=5/3$, and ④ $\mathit{A}\mathrm{Er}=-8<\mathcal{W}$, ⑤ $\mathit{A}\mathrm{Er}=\mathcal{W}\approx-3.55$, and ⑥ $\mathit{A}\mathrm{Er}=4>\mathcal{W}$. Color indicates the maximal shear stress, $\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$, within the cylinder and the blue curves are director lines. (bottom) The critical active Ericksen number, $\mathit{A}\mathrm{Er}=\mathcal{W}$, at which the deformed cylinder has aspect ratio one, plotted as a function of $w$ for $\kappa=5/3$. The points ①–③ correspond to the configurations shown in Fig. 2.
@@
The fluid flow triggered by the presence of a boundary may also be advantageous in biological cells, settings in which a disperse distribution of molecules, such as nutrients, chemicals, or oxygen, benefit the system’s function [75, 76, 77]. Consider an active nematic LC bounded inside, and tangentially anchored to, a cylinder with anchoring strength $w$. Here, the director angle of the LC can be written as
$\theta=\arg(1-\rho^{2}z^{2}),$ (21)
for the effective radius introduced in the external problem, $\rho(w)\coloneqq({\sqrt{1+4/w^{2}}-2/w})^{1/2}$, while the activity-induced flow can be expressed in terms of a streamfunction
$\begin{split}\psi=\frac{\mathit{A}}{8}(|z|^{2}-1)\operatorname{Im}\Big{[}&\frac{1-\rho z}{z^{2}}\log(1-\rho z)\\ +&\frac{1+\rho z}{z^{2}}\log(1+\rho z)\Big{]},\end{split}$ (22)
@@
see [66]. As before, the elasticity, activity, and viscosity of the LC will induce surface stresses on the boundary, which, if compliant, will deform. Figure 4(a) shows the director lines inside a cylinder that has deformed according to linear elasticity [66]. The final deformed shape resembles a tactoid [78], with a localized deformation at the left and right poles occurring for large anchoring strengths due to the appearance of two $+1$ aster defects when $w=\infty$. The deformed shape is dependent on the active Ericksen number, $\mathit{A}\mathrm{Er}$, with extensile activity ($A<0$) leading to parallel elongation and sufficient contractile activity ($A<0$) yielding perpendicular elongation [66], which is opposite to the aspect ratio’s depencend on the activity strength in the external problem.
@@
Figures 4(c–e) show snapshots of the fluid stirring induced by extensile activity. Varying the activity and anchoring strengths primarily affects the speed of stirring, though contractile activity also reverses its direction. The maximum fluid speed, $|\bm{u}|_{\max}=\max_{\bm{x}}\sqrt{u^{2}+v^{2}}$, is plotted as function of the anchoring strength in Fig. 4(b), for both the LC external and internal to a cylinder. The speeds plateau for large anchoring strengths, $w\gtrsim 10$, and decrease according to $|\bm{u}|_{\max}\sim 2|\mathit{A}|w/(25\sqrt{5})$ and $|\bm{u}|_{\max}\sim|\mathit{A}|w^{2}/(600\sqrt{5})$ as $w\to 0$, for the external and internal problem, respectively. The extent to which this stirring flow results in mixing is an interesting but separate question [79]. Mixing is notoriously challenging in highly viscous flows and in confined geometries [80], but is likely promoted by a wide spatial distribution of active stress [35, 36].
@@
Figure 4: (a) Deformation of a soft cylinder enclosing a passive nematic LC with $\mathit{A}\mathrm{Er}=0$, $w=10$, $M=50$, and $\kappa=5/3$. Color delineates the maximum shear stress, $\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$, within the solid and the blue curves are director lines. (b) Plot of the maximum fluid speed, $|\bm{u}|_{\max}=\max_{\bm{x}}\sqrt{u^{2}+v^{2}}$, as a function of $w$ for the LC external and internal to a cylinder. The asymptotic behaviour as $w\to 0$, i.e. $|\bm{u}|_{\max}\sim 2|\mathit{A}|w/(25\sqrt{5})$ (external) and $|\bm{u}|_{\max}\sim|\mathit{A}|w^{2}/(600\sqrt{5})$ (internal), are shown as black dashed lines. (c–e) Snapshots at times (c) $t=0$, (d) $t=100/|A|$, and (e) $t=200/|A|$ of colored fluid particles transported due to extensile activity inside the cylinder for $w=10$.
Conclusion:
We have shown that the configurations and dynamics of an active nematic can be affected dramatically by the mere presence of an immersed body, or boundary, for any non-trivial anchoring strength. The spontaneous flow that arises is characterized by plumes whose structure is anchoring strength dependent. The elastic, viscous, and active stresses may then deform the body, or external boundary, in a generally localized manner near (virtual) topological defects. Such flows can also produce a spontaneous stirring, and likely mixing, of the environment.
In the above theory, the deformation of the solid was assumed to be small enough that it does not affect the director field. In general, changes in the LC–solid interface could cause the director field to relax towards a different energy-minimizing configuration. Furthermore, although the flow does not directly influence the director field here, boundary deformations do depend on fluid flow, and so, this could provide an indirect influence if the interface deformation was taken into account: flows deform soft boundaries, thereby altering the LC. Even for a simple cylindrical domain, the deformation qualitatively depends on the activity strength (Fig. 3). Thus, since elongated colloids are known to reorient with the preferred axis of the surrounding LC and virtual defects seek out the corners of immersed bodies [70], significant changes to the LC could be expected here. The above work provides the first steps in modelling these complex fluid-structure interactions.
The localized deformations observed have also been reported in tactoids [78] and giant unilamellar vesicles [81]. Even for stiff bodies, the developed points could cause pinning of topological defects, which is important in setting the attraction angle of flexible colloid chains [82], for instance. Additionally, these deformations can encourage surface localized bundling of actin filaments [83, 84].
@@ Conclusion:
Looking to the horizon, biological cells are host to a multitude of active anisotropic networks. Can such fluids be triggered into motion solely by the passive anchoring conditions on an arriving organelle? This mechanism could offer an energy-efficient way to generate precisely timed mixing flows within the cellular environment.
@@ Acknowledgements.
Support for this research was provided by the Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation.

### v3

@@ III Surface tractions and soft boundaries
If a boundary is soft, it will generally be deformed by the surface tractions above. As a first step towards understanding more complex materials, we assume the boundary is part of an isotropic solid that deforms according to plane stress/strain in the framework of linear elasticity, with shear modulus $\mu_{s}$ and Poisson’s ratio $\nu$ [90, 91]. At equilibrium, the elastic solid can be described by an Airy-stress function, $\mathcal{A}(x,y)$, which satisfies the biharmonic equation, $\nabla^{4}\mathcal{A}=0$. We can, thus, write
@@
for two (new) Goursat functions, $G$ and $H$, that are locally holomorphic within the solid (but not necessarily the LC), analogous to $g$ and $h$ in (12) [90]. These Goursat functions yield the symmetric stress tensor within the solid, $\Sigma_{ij}$, via the Kolosov–Muskhelishvili formulae,
$\Sigma_{11}-\Sigma_{22}-2\mathrm{i}\Sigma_{12}=2\mathrm{i}\bar{z}G^{\prime\prime}(z)+2\mathrm{i}H^{\prime\prime}(z)$ (18a)
$\text{and}\quad\Sigma_{11}+\Sigma_{22}=4\operatorname{Im}G^{\prime}(z),$ (18b)
and the solid displacement, $(U,V)$, can be expressed as
@@
up to an arbitrary body displacement and rotation, for the dimensionless elastic modulus $\mathit{M}\coloneqq\mu_{s}L^{2}/K$ and $\kappa=(3-4\nu)$ or $(3-\nu)/(1+\nu)$, when assuming plane strain or plane stress, respectively.
Stress balance at the LC-solid interface, $\bm{\hat{\nu}}\cdot\bm{\Sigma}+\bm{t}^{e}+\bm{t}^{a}=\bm{0}$, yields the boundary conditions for $G$ and $H$:
$\partial_{s}\left[\bar{z}G^{\prime}(z)-\overline{G(z)}+H^{\prime}(z)\right]=(t^{e}_{x}+t^{a}_{x})-\mathrm{i}(t^{a}_{y}+t^{e}_{y}).$ (20)
III.1 Activity-controlled deformations
Returning to the example of a unit cylinder with tangential anchoring, the director angle (13) and streamfunction (14) yield analytic expressions for the elastic traction, $t^{e}_{x}+\mathrm{i}t_{y}^{e}$, and activity-induced traction, $t^{a}_{x}+\mathrm{i}t_{y}^{a}=\mathit{A}\mathrm{Er}(1/z-\rho^{2}z)$. The cylinder is now assumed to be deformable, and so, we seek the Airy stress (17). Here, the Goursat functions $G(z)$ and $H(z)$ must be locally holomorphic in $|z|<1$ and satisfy (20) on $|z|=1$; the solution is provided in [82].
Without activity ($\mathit{A}=0$), boundary deformations help to reduce the elastic energy stored in the bulk LC [80]. Figure 2 shows the LC configurations and body deformations for three different anchoring strengths and $\kappa=5/3$ (i.e. an incompressible cylinder, $\nu=1/2$, under plane stress). For large anchoring strengths, $w\gg 1$, most of the LC energy is stored local to the left and right poles of the cylinder, with a singular energy emerging in the case of strong (infinite) anchoring ($w=\infty$) due to the appearance of two $-1$ Boojum topological defects at $z=\pm 1$ [92, 88]. This localized energy within the LC results in large localized surface tractions as $w\to\infty$: $t^{e}_{x}+\mathrm{i}t_{y}^{e}=\mathcal{O}(w^{2})$ when $z\pm 1=\mathcal{O}(1/w)$, and $t^{e}_{x}+\mathrm{i}t_{y}^{e}=\mathcal{O}(1)$ otherwise. This, in turn, leads to large localized deformations; in particular, we find that
@@ III.1 Activity-controlled deformations
$\mathcal{U}(Z/w)=\frac{\kappa}{1+Z}+\frac{1}{1+\bar{Z}}+\frac{Z+\bar{Z}}{(1+\bar{Z})^{2}},$ (22)
corresponds to a deformation that is only apparent when $z\pm 1=\mathcal{O}(1/w)$. As the anchoring strength decreases, stress is no longer concentrated at the poles and the maximum shear stress $\tau(x,y)=\sqrt{(\Sigma_{11}-\Sigma_{22})^{2}/4+\Sigma_{12}^{2}}$ (a common indicator of mechanical failure [91]) is greatest at (±$(\pm x_{\max},0)$, where $x_{\max}$ decreases from $1$ when $w=\infty$, to 00 when $w\leq 24/35$, as shown in Fig. 2.
Figure 2: (top) Deformation of a soft cylinder immersed in a passive nematic LC with $\mathit{A}\mathrm{Er}=0$, $\mathit{M}/w=5$, $\kappa=5/3$, and dimensionless anchoring strengths ① $w=0.1$, ② $w=1$, and ③ $w=10$. Color indicates the maximum shear stress, $\tau=\sqrt{(\Sigma_{11}-\Sigma_{22})^{2}/4+\Sigma_{12}^{2}}$, within the cylinder and the blue curves are director lines. (bottom) The positions, $(x_{\max},0)$, and value, $\tau_{\max}=\max_{\bm{x}}\tau$, of the largest maximum shear stress plotted as a function of $w$.
In an active nematic ($A\neq 0$), the activity leads to an additional deformation:
$U+\mathrm{i}V=U^{e}+\mathrm{i}V^{e}+\frac{\mathit{A}\mathrm{Er}}{4\mathit{M}}\bigl{[}(1-\kappa)\rho^{2}z+2\bar{z}\bigr{]},$ (23)
@@
$\mathcal{A}_{R}\sim 1+U\bigr{|}_{z=1}-V\bigr{|}_{z=\mathrm{i}}=1+\frac{\mathit{A}\mathrm{Er}-\mathcal{W}}{\mathit{M}},$ (24)
@@
Figure 3: (top) Deformation of a soft cylinder immersed in an active nematic LC with $w=5$, $\mathit{M}=50$, $\kappa=5/3$, and ④ $\mathit{A}\mathrm{Er}=-8<\mathcal{W}$ (extensile), ⑤ $\mathit{A}\mathrm{Er}=\mathcal{W}\approx-3.55$ (extensile), and ⑥ $\mathit{A}\mathrm{Er}=4>\mathcal{W}$ (contractile). Color indicates the maximal shear stress, $\sqrt{(\Sigma_{11}-\Sigma_{22})^{2}/4+\Sigma_{12}^{2}}$, within the cylinder and the blue curves are director lines. (bottom) The critical active Ericksen number, $\mathit{A}\mathrm{Er}=\mathcal{W}$, at which the deformed cylinder has aspect ratio one, plotted as a function of $w$ for $\kappa=5/3$. The points ①–③ correspond to the configurations shown in Fig. 2.
III.2 Bounded domains and spontaneous stirring
@@
The fluid flow triggered by the presence of a boundary may also be advantageous in interior domains like biological cells — settings in which a distribution of molecules, such as nutrients, chemicals, or oxygen, benefit the system’s function [93, 94, 95, 96]. Consider an active nematic LC bounded inside and (weakly) tangentially anchored to a cylinder with anchoring strength $w$. Here, the director angle of the LC can be written as $\theta=\arg(1-\rho^{2}z^{2})$, for the effective radius introduced in the external problem, $\rho(w)=({\sqrt{1+4/w^{2}}-2/w})^{1/2}$, while the flow can be expressed in terms of a streamfunction (see [82]):
$\begin{split}\psi=\frac{\mathit{A}}{8}(|z|^{2}-1)\operatorname{Im}\Big{[}\frac{1-\rho z}{z^{2}}&\log(1-\rho z)\\ +&\frac{1+\rho z}{z^{2}}\log(1+\rho z)\Big{]}.\end{split}$ (26)
@@
Figure 4(a) shows the director lines inside a cylinder that has deformed according to linear elasticity [82]. The final deformed shape resembles a tactoid [97], with a localized deformation at the left and right poles occurring for large anchoring strengths due to the appearance of two $+1$ aster defects when $w=\infty$. The deformed shape is dependent on the active Ericksen number, $\mathit{A}\mathrm{Er}$, with extensile activity ($A<0$) leading to parallel elongation and sufficient contractile activity ($A<0$) yielding perpendicular elongation [82] — opposite to the dependence on the activity strength in the external problem.
@@
Figures 4(c–e) show stirring induced by extensile activity. Varying the activity and anchoring strength affects the rate of stirring. The maximum fluid speed, $|\bm{u}|_{\max}$, is plotted as function of anchoring strength in Fig. 4(b), for both the external and internal problems. The speeds plateau for large anchoring strengths, and decrease according to $|\bm{u}|_{\max}\sim 2|\mathit{A}|w/(25\sqrt{5})$ and $|\bm{u}|_{\max}\sim|\mathit{A}|w^{2}/(600\sqrt{5})$ as $w\to 0$, for the external and internal problem, respectively. The extent to which this stirring flow is mixing is an interesting but separate question [98]. Mixing is notoriously challenging in highly viscous flows and in confined geometries [99], but is likely promoted by a wide spatial distribution of active stress [48, 49].
@@
Figure 4: (a) Deformation of a soft cylinder enclosing a passive nematic LC with $\mathit{A}\mathrm{Er}=0$, $w=10$, $M=50$, and $\kappa=5/3$. Color delineates the maximum shear stress, $\sqrt{(\Sigma_{11}-\Sigma_{22})^{2}/4+\Sigma_{12}^{2}}$, within the solid and the blue curves are director lines. (b) Plot of the maximum fluid speed, $|\bm{u}|_{\max}=\max_{\bm{x}}\sqrt{u^{2}+v^{2}}$, as a function of $w$ for the LC external and internal to a cylinder. The asymptotic behaviours as $w\to 0$ are shown as black dashed lines. (c–e) Snapshots at times (c) $t=0$, (d) $t=100/|A|$, and (e) $t=200/|A|$ of colored fluid particles transported due to extensile activity inside the cylinder for $w=10$.
IV Conclusions and discussions
We have shown that the configurations and dynamics of an active nematic can be affected dramatically by the mere presence of a boundary, for any non-trivial anchoring strength. It suggests that energy bound in a uniform active nematic can be released as a large scale flow due to a localized inclusion alone.
@@ IV Conclusions and discussions
The spontaneous flow that arises is characterized by unique plumes whose structure is anchoring strength dependent. The body may then deform in response to this anisotropic, viscoelastic flow, in a generally localized manner near (virtual) topological defects. Such flows can also produce a spontaneous stirring, and likely mixing, of the environment. The associated symmetry breaking can lead to cell or droplet division and motility [100, 101, 102]. Looking to the horizon, biological cells are host to a multitude of active anisotropic networks. Can such fluids be triggered into motion solely by the passive anchoring conditions on an arriving organelle? This mechanism could offer an energy-efficient way to generate precisely timed mixing flows within the cellular environment.
New experiments on the shapes and dynamics of deformable inclusions in active nematics is needed. Testable hypotheses include the differential direction of elongation, which may be tuned by varying either the activity strength (e.g. by changing the density of molecular motors  [23]) or by changing the body stiffness. With sufficient perpendicular elongation due to strong extensile activity, as in the top left panel of Fig. 3, body rotation may also ensue to reduce the bulk elastic energy. We thus predict that a diverse zoology of body dynamics will emerge with increasing extensile activity in particular.
[…]
Future directions will require different machinery. Large deformations will more completely couple the LC configuration and flows, and a finite rotational viscosity will introduce an additional timescale [77, 109]. The analytical limits provided herein may still be of use for steering investigations of these more highly nonlinear regimes.
@@ Acknowledgements.
We acknowledge helpful discussions with Nicholas L. Abbott and Ido Lavi. Support for this research was provided by the Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation.